A bug on the surface of a pond is observed to move up and down atotal vertical distance of 5.0 cm,lowest to highest point, as a wave passes. If the ripples decreaseto 4.7 cm, by what factor does thebug's maximum KE change?
Answer #1

We know the formula for the kinetic energy of the particlewhich executes the simple harmonic motion is

$$ \mathrm{K}=(1 / 2) \mathrm{m} \mathrm{A}^{2} \omega^{2} $$

From this formula kinetic energy \((\mathrm{K})\) is directly proportionalto square of the amplitude.

Then

$$ \frac{K_{2}}{K_{1}}=\frac{A_{2}^{2}}{A_{1}^{2}} $$

$$ =\frac{(4.7 \mathrm{~cm})^{2}}{(5.0 \mathrm{em})^{2}}=0.8836 $$